OpenStudy (anonymous):

Simplify the expression: Simplify the expression: (6x^3 − 7x − 4) - (4x^3 + 8x + 3). ^ indicates exponent

OpenStudy (anonymous):

combine like terms.

OpenStudy (anonymous):

like terms are terms where literal factors are the same, example: 3x^3, and 1x^3 are like terms.

OpenStudy (anonymous):

again same?

OpenStudy (anonymous):

ok, I got 2x^3+x+1 and no, it's not the same.

OpenStudy (anonymous):

Let me give you an example on combining like terms: 5x^3 - 2x^3 = 3x^3

OpenStudy (anonymous):

check the +x +1. part

OpenStudy (anonymous):

yes just subtraction here.

OpenStudy (anonymous):

-7x - 8x = ___? example: -2x - 3x = -5x.

OpenStudy (anonymous):

ok, so the last term is a negatve? = 2x^3+x-1

OpenStudy (anonymous):

sorry, not quite.

OpenStudy (anonymous):

can you draw something out for me so I can see where I'm going wrong please? :)

OpenStudy (anonymous):

okay,